OpenStudy (highschoolmom2010):

A class has 30 students. In how many ways can committees be forced using the following numbers of students? a. 3 students b. 5 students

OpenStudy (raden):

these is the combination problem. just use the formula : nCr = n!/(n-r)!r! in this case n = 30 part a, given r = 3 so, it is 30C3 = 30!/(30-3)3! = 30!/27!3! = 30 * 29 * 28 * 27!/(27! * 3 * 2 * 1) = 4060 ways. now can you get part b ?

OpenStudy (highschoolmom2010):

im writing the part a one second :))

OpenStudy (cwrw238):

another 'shortcut' way to work these out is for example for a@ number if combnations = 30 * 29 * 28 ---------- 3 * 2 * 1

OpenStudy (cwrw238):

b would be 30 * 29 *28*27*26 --------------- 5 *4 * 3 * 2 * 1

OpenStudy (highschoolmom2010):

im lost on part a how you got 4060

OpenStudy (raden):

30 * 29 * 28 * 27!/(27! * 3 * 2 * 1) cancel the 27!'s, get 30 * 29 * 28 / 3 * 2 * 1 = 24360/6 = 4060 part b. just put r = 5 into that formula. mom :)

OpenStudy (cwrw238):

that's basically what I did - I just cancelled the 27! in my head in part b - you cancel the 25!

OpenStudy (cwrw238):

do you see the pattern in my calculation? you can then cancel out the numbers further to make the calculation easier: